# Quantum anomalous Hall phases in gated rhombohedral graphene

### v1

@@ 2 Effective model and classification
Note that for $0\leq j\leq m$ we have that:
$i\lim_{k\to 0}\big{(}\psi_{j}^{\pm}(k),U^{\prime}(0)\psi_{j}^{\pm}(k)\big{)}=j.$
Also note from (11) that phase transitions happen when $j-\frac{m}{2}=\delta(\gamma,u)$. First consider $\delta_{0}\leq\frac{1}{2}<\delta(\gamma,u)<\delta_{1}$ with $m$ even for simplicity. For $m$ even, $\delta_{1}=1$ so that $E_{j}^{\pm}>0$ for $0\leq j\leq\frac{m}{2}-1$ and $E_{m/2}^{+}>0$. Therefore, summing over positive eigenvalues for $u<0$ gives:
$i\lim_{k\to 0}\sum_{E_{j}^{\pm}>0}\big{(}\psi_{j}^{\pm}(k),U^{\prime}(0)\psi_{j}^{\pm}(k)\big{)}=2\sum_{j=1}^{\frac{m}{2}-1}j+\frac{m}{2}=\frac{m}{2}\big{(}\frac{m}{2}-1\big{)}+\frac{m}{2}=\frac{m^{2}}{4}.$
Similarly for $u>0$ we get:
$i\lim_{k\to 0}\sum_{E_{j}^{\pm}>0}\big{(}\psi_{j}^{\pm}(k),U^{\prime}(0)\psi_{j}^{\pm}(k)\big{)}=\frac{m}{2}+2\Big{(}\sum_{j=\frac{m}{2}+1}^{m-1}j\Big{)}+m=\frac{3m^{2}}{4}$
so that ${\rm BDI}_{0}=\frac{m^{2}}{2}$. Now notice that if $\delta_{j}<\delta(\gamma,u)<\delta_{j+1}$ then $E_{\frac{m}{2}-k}^{-}<0$ and $E_{\frac{m}{2}+k}^{+}>0$ for $1\leq k\leq j$. Therefore ${\rm BDI}_{j}$ is given by exchanging the associated curvature terms for the eigenvalues which change signs from $\mathcal{C}[\Pi^{S}]$ to $\mathcal{C}[\Pi^{N}]$ and vice versa:
$\begin{array}[]{rcl}{\rm BDI}_{j}&=&{\rm BDI}_{0}+\displaystyle\sum_{l=1}^{j}\left[\left[\left(\frac{m}{2}-l\right)-\left(\frac{m}{2}+l\right)\right]-\left[\left(\frac{m}{2}+l\right)-\left(\frac{m}{2}-l\right)\right]\right]\\[8.53581pt] &=&{\rm BDI}_{0}-4\displaystyle\sum_{l=1}^{j}l\ =\ \frac{m^{2}}{2}-2\delta_{j}(\delta_{j}+1).\end{array}$
A similar calculation for $m$ odd yields:
$\begin{array}[]{rcl}{\rm BDI}_{j}&=&\dfrac{m^{2}}{2}+\dfrac{1}{2}+\displaystyle\sum_{l=0}^{j}\left[\left[\left(\lfloor\frac{m}{2}\rfloor-l\right)-\left(\lceil\frac{m}{2}\rceil+l\right)\right]-\left[\left(\lceil\frac{m}{2}\rceil+l\right)-\left(\lfloor\frac{m}{2}\rfloor-l\right)\right]\right]\\ &=&\dfrac{m^{2}}{2}+\dfrac{1}{2}-2\displaystyle\sum_{l=0}^{j}(2l+1)=\frac{m^{2}}{2}-2(j^{2}+2j+\frac{3}{4})=\frac{m^{2}}{2}-2\delta_{j}(\delta_{j}+1).\end{array}$
@@ 4 Numerical simulations
We start with cross sections of the rotationally invariant bulk spectrum of the 5-layer system (4) (involving diagonalization of $10\times 10$ matrices performed in Matlab) in Figure 2 for various values of $u$ assuming a fixed coupling constant normalized to $\gamma=1$. For the low-potential (relative to $\gamma$) case we recover the band structure noted in [31] for tri-layers which exhibits two minima (maxima) in the bands closest to $E=0$. In the high-potential case, we recover the pattern discussed in [7] where for $m$ layers, $m$ minima (maxima) are observed in the bands closest to $E=0$ and the distance from $E=0$ scales as $(\gamma/u)^{2l}$ for $l=\{1,...,\lceil m/2\rceil\}$. For intermediate values of $u$, we observe an intermediate number of minima close to $E=0$. Importantly, the size of the global band gap scales proportional to $(\gamma/u)^{2m}$ for $u$ sufficiently large as described in [7] so that the global gap can be arbitrarily small as $u,m\to\infty$, even if they are guaranteed to remain positive by Theorem 2.2. This is illustrated in Figure 2(d) where the gap is barely discernible for the highest $|\mathbf{k}|$ at which a minimum occurs.
The theory in [7] ensures that the smallest spectral gap appears at the large values of $|k|$, which is somewhat reassuring given that high-wavenumber oscillations may be able to be disregarded in real systems. However, for some parameter regimes close to critical values, we find that arbitrarily small spectral gaps can also be found for a range of $|k|$ values close to 0. In Figure 3 we demonstrate this for $m=9$, where a spectral gap of $\approx 10^{-3}$ is seen at $|k|\approx\pm 0.32$, creating an obstacle to observing edge states experimentally. This finding may also call into question whether the topological properties of the system are robust to perturbations due to the fact that our analysis relies heavily on a global gap which could be closed by exceedingly small perturbative effects at these parameter values.
Figure 2: Bulk spectrum for 5-layer model with $\gamma=1$ and $u=(1,2.5,3.5,7)$ from left to right, top to bottom. For the low-potential case we recover the “Mexican hat” shape described in [23, 32] for the bands nearest $E=0$ and for the high-potential case we note 5 minima (maxima) in the positive (negative) band closest to $E=0$ as described in [7]. Intermediate values of $u$ produce an intermediate number of minima.
Figure 3: Bulk spectrum for $m=9$ layers. As number of layers increases the global band gap near phase transitions becomes increasingly small even for intermediate values of $k_{x}$. Middle frame shows the branch closest to $E=0$ and right from shows detail of the minimum in said branch closest to $k_{x}=0$. Spectrum shown is for $(\gamma,u)=(1,2.25)$
@@
In order to calculate the spectrum of interface Hamiltonians numerically, we assume that $H$ is invariant in the x-direction and $u(y)$ varies from positive to negative as a function of $y$ as described above. Taking the Fourier transform in $(x,t)\to(k_{x},E)$ under these assumptions gives an interface Hamiltonian:
@@
Assuming that for large enough $|y|$, the coefficients of $A(k_{x},E)$ are constant and given by the constant-coefficient operator $A^{N/S}(k_{x},E)$ for $y\geq R$ and $y\leq-R$ respectively, then for any fixed $(k_{x},E)$ we can determine the bulk modes by simply diagonalizing $A^{N},A^{S}$. In general, however, an explicit solution for $(\psi(y),E)$ is not feasible analytically for the ODE (20). A numerical ODE solver is used to solve (20) for $\psi(y=0)$ by using the eigenvectors of $A^{N}$ and $A^{S}$ as initial conditions at $y=\pm R$ respectively and numerically solving for $\psi(y)$ for R<y<0-R<y<0 and $0<y<R$ respectively. Exponentially increasing eigenvectors from the bulk spectrum of $A^{N}$ and $A^{S}$ are eliminated as non-physical. If the subspaces of valid (non-exponentially increasing) eigenvectors of $A^{N},A^{S}$ intersect non-trivially once evaluated at $y=\pm 0$ then there exists a valid solution of (20). More precisely, orthogonal projectors $\Pi^{N},\Pi^{S}$ onto the subspaces of valid eigenvectors of $A^{N},A^{S}$ evaluated at $y=0$ are formed, and a $(k_{x},E)$ pair is accepted if the largest eigenvalue of $\Pi^{N}\Pi^{S}$ is within a small tolerance of 1 ($10^{-4}$ for Figures 3-6). When compared with widely-used finite difference methods [13, 8, 7] this method has no need for periodization of the domain and therefore does not require the heuristic elimination of modes [8, 7] or alternatively the necessity of modeling two equal but opposite transitions [13, 31]. Comparison with finite-difference methods (not displayed here) does however show that the two methods agree using a fine enough mesh with finite differences.
Figure 4 validates the values of the BDI given in Theorem 2.3 for the cases of $3\leq m\leq 6$ layers and for all possible transition values of $\delta_{j}$. As shown in Theorem 2.3 and Fig. 1, $m=3$ demonstrates only one distinct topological phase. For $m=4,5,6$ we derive from (13) the critical values $u_{c}=2\sqrt{2}\gamma$ ($m=5$), $u_{c}=2\sqrt{3}\gamma$ ($m=4$), and $u_{c}=(10\gamma/\sqrt{3},10\gamma/\sqrt{15})$ ($m=6$). Figure 4 contains numerically calculated spectra for all phases divided by these critical values assuming $\gamma=1$. We also validate in Figure 5 using 5 layers that band crossings only occur at $k=0$ and show the transition from 5 to 11 edge modes at the critical value $u=2\sqrt{2}\gamma$. All results are in perfect agreement with the results of Theorems 2.2 and 2.3.
Figure 4: Numerically calculated spectra for $m=\{3,4,5,6\}$ for all phases. We assume $\gamma=1$ for simplicity and calculate the spectrum of $H_{I}$ for $u=2$ $(m=3)$, u = (5, 2) ($m=4$), $u=(4,1.2)$ ($m=5$), and $u=(6,4,2)$ ($m=6$). The approximate critical values from (13) are $u_{c}\approx 3.46$ (m = 4), $u_{c}\approx 2.83$ (m = 5), and $u_{c}\approx(2.58,5.77)$ ($m=6$). The number of edge states agrees with Figure 1 for $m=3,4,5,6$ and all respective phases.
@@
Figure 6: Eigenvectors of edge states in the 3-layer case ($(u,\gamma)=(2,1)$) at $E=0$. Left is the eigenvector as a function of $y$ at $(k_{x},E)=(0,0)$ and right at $(k_{x},E)=(-1,0)$. The edge is located within 1<y<1-1<y<1. $A_{n}$ and $B_{n}$ labels correspond to the $A$ and $B$ sub-lattice components of layer $n$ following the notation of [32]. Polarization is seen biased toward the outer layers.
Note that the edge of the band gap always appears flat between two corresponding peaks of the bulk spectrum. This is a consequence of rotational invariance of the Hamiltonian. In Figure 2 the bulk spectrum is plotted as a function of $|\mathbf{k}|$. However, by rotational invariance the energy at $\mathbf{k}=(k,0)$ is equivalent to the energy at $\mathbf{k}=(k_{x},k_{y})$ given $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$. Therefore each point $(k_{x},E)$ in Figure 2 can be mapped to a point equal in energy and closer to $k_{x}=0$ by adding an appropriate $k_{y}$ value; e.g. $(k_{x},E)\to(k_{x}^{\prime},E)$ where $k_{x}=\sqrt{k_{x}^{{}^{\prime}2}+k_{y}^{2}}=|\mathbf{k}|$.
@@ 5 Discussion
The result of Theorem 2.3 gives all possible topological values of the quantum anomalous Hall effect in an idealized macroscopic (effective) model of multi-layer rhombohedral graphene and Floquet topological insulators. In the latter application, we retrieve the topological invariant obtained for large values of the driving laser frequency (corresponding to a small coupling constant $\gamma$) [7]. In the former application, we retrieve the topological invariant obtained for small displacement field [17, 31] compared to the coupling constant $\gamma$.
@@
We also note that instead of modeling a transition $u\to-u$ we could consider the case that across a large enough length scale the energetically preferable stacking order changes continuously from AB to BA, as considered in [6]. This transition is modeled by a continuous transition of $B\to B^{*}$ instead of $u\to-u$. Substituting the eigenvectors of $\hat{H}(\gamma A^{*})$ for $\hat{H}(-u)$ in Section 3 using the unitary relations derived in (8) it is clear that the BDI’s for this transition are identical to the ones derived in Section 3. Figure 7 shows for the 5-layer case that indeed these two transitions are topologically equivalent.
In the RHG application, we assume here a model (1) that is both spin- and valley-polarized. While the obtained numbers do not depend on spin, they depend on the valley index $\tau=\pm 1$. The total QAH therefore vanishes unless valley polarization may be obtained experimentally, which is achieved in [17] by means of an appropriate spin orbit coupling. It is unclear whether such polarization may still be achieved for the large displacement fields (large values of $u$) necessary to observe a phase transition (i.e., $\delta<\delta_{\lfloor\frac{m}{2}-1\rfloor}$).
The theoretical results, and in particular the presence of a gap at $E=0$, depend on the specific structures of $H$ in (1). More general interlayer couplings such as those proposed in [17, 31] do not modify the elliptic nature of the operators and the bulk-edge correspondence of [4, 26] still applies. Therefore, any continuous deformation of the model in (1) with bulk phases that does no close the gap at $E=0$ will generate an edge asymmetry for which the results and explicit indices of Theorem 2.3 apply. While the results of Theorem 2.2 guarantee that gap crossings may only occur at $E=0$, we found examples of values of $(\gamma,u)$ such that the spectral gap at values $k\not=0$ may be significantly smaller than the gap at $k=0$; see for instance Fig. 3. When considering interactions between more distant layers, as in e.g. [32], rotational symmetry which has been essential to our analysis here breaks down and the perturbative effects of further off-diagonal elements have yet to be explored. It is therefore unclear whether such gaps persist except for very small perturbations of the model (1).
@@
Figure 7: Edge spectrum for 5-layer case with $A\to A^{*}$ transition. Parameter values used are $u=(4,1.2)$ and $\gamma=1$. Comparison with Figure 4 shows that the number of edge states agrees for both transitions.
@@
In summary, we have found all the topological phases and bulk difference invariants of an effective model of coupled 2-d Dirac systems which has applications in gated ABC-stacked rhombohedral graphene with an arbitrary number of layers [32, 23, 31] and Floquet topological insulators [7]. Due to the ellipticity of the problem the bulk-edge correspondence holds [4, 3] and we verify by numerical spectral calculations the number of predicted edge modes for all topological phases in the $m=\{3,4,5,6\}$ cases. For $u<\gamma$ we recover the results of [31] for the QAH effect in RHG and for $u\gg\gamma$ the results of [7] for Floquet topological insulators. Our results show that for an $m$-layer model there are in fact $\lfloor\frac{m}{2}\rfloor$ topological phases in between the aforementioned parameter regimes whose BDI is given by (14). Although it remains to be seen if these phases are robust to perturbations in interactions beyond the nearest-layer interactions we consider here, these results show the exciting possibility of tunable QAH states in multi-layer ABC stacked graphene should experimental conditions allow high enough gate potentials to be achieved.
Appendix A Appendix

### v2

@@ 2 Effective model and classification
Note that for $0\leq n\leq m$ we have that:
$i\lim_{k\to 0}\big(\psi_{n}^{\pm}(k,0),U^{\prime}(0)\psi_{n}^{\pm}(k,0)\big)=n.$
Computation of BDI:
Recall from (14) that $\hat{H}_{B}(\mathbf{k},\gamma,u)$ is in phase $j$ when $\delta_{j-1}<\delta(\gamma,u)<\delta_{j}$. Denote $\Pi^{(j)}$ by (9) with $\hat{H}_{B}(\mathbf{k})$ in phase $j$. First, we compute $\mathcal{C}[\Pi^{(\pm 1)}]$ for $m$ even, then extend to arbitrary phase $j$ and odd $m$. For $m$ even $\delta_{1}=1$ so that in phase 1 $E_{n}^{\pm}<0$ for $0\leq n\leq\frac{m}{2}-1$ and $E_{m/2}^{-}<0$ since $0<\delta(\gamma,u)<1$. Therefore we obtain:
$\mathcal{C}[\Pi^{(1)}]=-i\lim_{k\to 0}\sum_{E_{n}^{\pm}<0}\big(\psi_{n}^{\pm}(k,0),U^{\prime}(0)\psi_{n}^{\pm}(k,0)\big)=2\sum_{n=1}^{\frac{m}{2}-1}n+\frac{m}{2}=\frac{m}{2}\big(\frac{m}{2}-1\big)+\frac{m}{2}=\frac{m^{2}}{4}.$
For phase -1 the results are identical except that $u<0$ so that $E_{n}^{\pm}<0$ for $\frac{m}{2}+1\leq n\leq m$ and $E_{\frac{m}{2}}^{+}<0$ which gives:
$\mathcal{C}[\Pi^{(-1)}]=-i\lim_{k\to 0}\sum_{E_{n}^{\pm}<0}\big(\psi_{n}^{\pm}(k,0),U^{\prime}(0)\psi_{n}^{\pm}(k,0)\big)=\frac{m}{2}+2\Big(\sum_{n=\frac{m}{2}+1}^{m-1}n\Big)+m=\frac{3m^{2}}{4}.$
In order to enforce the symmetry $\mathcal{C}[\Pi^{(j)}]=-\mathcal{C}[\Pi^{(-j)}]$ (which we may do WLOG since our final result involves differences of $\mathcal{C}[\Pi^{(j)}]$’s) we apply another global gauge transformation of $\psi_{j}(k,\theta)\to e^{-i\frac{m^{2}}{4}\theta}\psi_{j}(k,\theta)$ so that:
$\mathcal{C}[\Pi^{(\pm 1)}]=\mp\frac{m^{2}}{4}.$
Now notice that if $\delta_{j-1}<\delta(\gamma,u)<\delta_{j}$ then for $u>0$, $E_{\frac{m}{2}+l}^{-}<0$ and $E_{\frac{m}{2}-l}^{+}>0$ for $1\leq l\leq j-1$. Therefore for $j>1$ we can obtain $\mathcal{C}[\Pi^{(j)}]$ by subtracting the $E^{+}_{\frac{m}{2}-l}$ and adding the $E_{\frac{m}{2}+l}^{-}$ curvature terms from $\mathcal{C}[\Pi^{(1)}]$ for $1\leq l\leq j-1$:
$\mathcal{C}[\Pi^{(j)}]=\mathcal{C}[\Pi^{(1)}]+\sum_{l=1}^{j-1}\left(\left(\frac{m}{2}+l\right)-\left(\frac{m}{2}-l\right)\right)=-\frac{m^{2}}{4}+2\sum_{l=1}^{j-1}l=-\frac{m^{2}}{4}+\delta_{j}(\delta_{j}-1).$
Similarly for $u<0$ we get $E_{\frac{m}{2}+l}^{-}>0$ and $E_{\frac{m}{2}-l}^{+}<0$ for $1\leq l\leq j-1$ so that for $j<1$:
$\mathcal{C}[\Pi^{(j)}]=\mathcal{C}[\Pi^{(-1)}]-\sum_{l=1}^{j-1}\left(\left(\frac{m}{2}+l\right)-\left(\frac{m}{2}-l\right)\right)=\frac{m^{2}}{4}-\delta_{j}(\delta_{j}-1)=-\mathcal{C}[\Pi^{(-j)}].$
Performing similar calculations with $m$ odd we find, with a global gauge transformation of $e^{-i(\frac{m^{2}}{2}-m)\theta}$, that:
$\mathcal{C}[\Pi^{(1)}]=-\mathcal{C}[\Pi^{(-1)}]=-\frac{m^{2}}{4}+\frac{3}{4}=-\frac{m^{2}}{4}+\delta_{1}(\delta_{1}-1)$
and subsequently for $j>1$:
$\mathcal{C}[\Pi^{(j)}]=-\mathcal{C}[\Pi^{(-j)}]=-\frac{m^{2}}{4}+\frac{3}{4}+\sum_{l=1}^{j-1}\left(\left((\lceil\frac{m}{2}\rceil+l\right)-\left(\lfloor\frac{m}{2}\rfloor-l\right)\right)\qquad\qquad$
$\qquad=-\frac{m^{2}}{4}+(j^{2}-\frac{1}{4})=-\frac{m^{2}}{4}+\delta_{j}(\delta_{j}-1).$
Thus for all $m$ and all phases $j\in\{\pm 1,\pm 2,...,\pm\lfloor\frac{m}{2}\rfloor\}$ we get:
$\mathcal{C}[\Pi^{(j)}]=-\text{sgn}(j)\left(\frac{m^{2}}{4}-\delta_{j}(\delta_{j}-1)\right)$ (21)
[…]
@@ 4 Numerical simulations
We start with cross sections of the rotationally invariant bulk spectrum of the 5-layer system (4) (involving diagonalization of $10\times 10$ matrices performed in MATLAB) in Figure 2 for various values of $u$ assuming a fixed coupling constant normalized to $\gamma=1$. For the low-potential (relative to $\gamma$) case we recover the band structure noted in [36] for tri-layers which exhibits two minima (maxima) in the bands closest to $E=0$. In the high-potential case, we recover the pattern discussed in [3] where for $m$ layers, $m$ minima (maxima) are observed in the bands closest to $E=0$ and the distance from $E=0$ scales as $(\gamma/u)^{2l}$ for $l=\{1,...,\lceil m/2\rceil\}$. For intermediate values of $u$, we observe an intermediate number of minima close to $E=0$ which varies from 2 to $m$ as $u$ increases. Importantly, the size of the global band gap scales proportional to $(\gamma/u)^{2m}$ for $u$ sufficiently large as described in [3] so that the global gap can be arbitrarily small as $u,m\to\infty$, even if the gap is guaranteed to exist by Theorem 2.2. This is illustrated in Figure 2(d) where the gap is barely discernible for the highest $|\mathbf{k}|$ at which a minimum occurs.
The theory in [3] ensures that the smallest spectral gap appears at the large values of $|\mathbf{k}|$, which is somewhat reassuring given that high-wavenumber oscillations may be able to be disregarded in real systems. However, for some parameter regimes close to critical values, we find that arbitrarily small spectral gaps can also be found for a range of intermediate $|\mathbf{k}|$ which are not too large nor exceedingly close to $|\mathbf{k}|=0$, where we know that band crossings occur according to (13). In Figure 3 we demonstrate this for $m=9$, where a spectral gap of $\approx 10^{-3}$ is seen at $|\mathbf{k}|\approx\pm 0.32$, creating an obstacle to observing edge states experimentally for higher values of $m$. This finding may also call into question whether the topological properties of the system are robust to perturbations due to the fact that our analysis relies heavily on a global gap which could be closed by exceedingly small perturbative effects at these parameter values.
Figure 2: Bulk spectrum for 5-layer model with $\gamma=1$ and $u=(1,2.5,3.5,7)$ for (a), (b), (c), and (d) respectively. For the low-potential case we recover the “Mexican hat” shape [27, 38] for the bands nearest $E=0$ and for the high-potential case we note 5 minima (maxima) in the positive (negative) band closest to E=0E=0[3]. Intermediate values of $u$ produce an intermediate number of minima.
Figure 3: (a) Bulk spectrum for $m=9$ layers. As number of layers increases the global band gap near phase transitions becomes increasingly small even for intermediate values of $|\mathbf{k}|$. (b) shows the branch closest to $E=0$ and (c) shows detail of the minimum in said branch closest to $k_{x}=0$. Spectrum shown is for $(\gamma,u)=(1,2.25)$.
@@
In order to calculate the spectrum of interface Hamiltonians numerically, we assume that $H$ is invariant in the x-direction and $u(y)$ varies between $u^{S}$ and $u^{N}$ in the region $R\leq y\leq R$ as described above. Taking the Fourier transform in $(x,t)\to(k_{x},E)$ under these assumptions gives an interface Hamiltonian:
@@
Assuming that for large enough $|y|$, the coefficients of $A(k_{x},E)$ are constant and given by the constant-coefficient operator $A^{N/S}(k_{x},E)$ for $y\geq R$ and $y\leq-R$ respectively, then for any fixed $(k_{x},E)$ we can determine the bulk modes by simply diagonalizing $A^{N},A^{S}$. In general, however, an explicit solution for $(\psi(y),E)$ is not feasible analytically for the ODE (22). A numerical ODE solver is used to solve (22) for $\psi(y=0)$ by using the eigenvectors of $A^{N}$ and $A^{S}$ as initial conditions at $y=\pm R$ respectively and numerically solving for $\psi(y)$ for R<y<0-R<y<0 and $0<y<R$ respectively. Exponentially increasing eigenvectors from the bulk spectrum of $A^{N}$ and $A^{S}$ are eliminated as non-physical. If the subspaces of valid (non-exponentially increasing) eigenvectors of $A^{N},A^{S}$ intersect non-trivially once evaluated at $y=\pm 0$ then there exists a valid solution of (22). More precisely, orthogonal projectors $P^{N},P^{S}$ onto the subspaces of valid eigenvectors of $A^{N},A^{S}$ evaluated at $y=0$ are formed, and a $(k_{x},E)$ pair is accepted if the largest eigenvalue of $P^{N}P^{S}$ is within a small tolerance of 1 ($10^{-4}$ for Figures 3-6). When compared with widely-used finite difference methods [17, 4, 3] this method has no need for periodization of the domain and therefore does not require the heuristic elimination of modes [4, 3] or alternatively the necessity of modeling two equal but opposite transitions [17, 36]. This method shares these advantages with two previously-proposed approaches[13, 12] but may also be applied to continuum models as well as periodic lattice models. Comparison with finite-difference methods (not displayed here) does however show that the two methods agree using a fine enough mesh with finite differences where spurious modes are eliminated.
[…]
Figure 4: Numerically calculated spectra for $m=\{3,4,5,6\}$ for symmetric transitions between the two phases j-j and $j$ for $j\in\{1,2,3\}$. We assume $\gamma=1$ for simplicity and calculate the spectrum of $H_{I}$ for $u=2$ $(m=3)$, u = (5, 2) ($m=4$), $u=(4,1.2)$ ($m=5$), and $u=(6,4,2)$ ($m=6$). The approximate critical values from (13) are $u_{c}\approx 3.46$ (m = 4), $u_{c}\approx 2.83$ (m = 5), and $u_{c}\approx(2.58,5.77)$ ($m=6$). Note that BDI’s for symmetric transitions are denoted by in Figure 1. The number of edge states agrees with Figure 1 for $m=3,4,5,6$ and all respective phases.
Figure 5: Illustration of the asymmetric transitions in $u$ for $m=5$ layers. (a) shows a transition from phase +1 to +2 where $u$ varies from $u^{S}=2$ to $u^{N}=4$. (b) shows a transition from phase -1 to phase 2 where $u$ varies from $u^{S}=-2$ to $u^{N}=4$. Both transitions are for $\gamma=1$, giving a critical value of $u_{c}=2.83$ as in the previous figure.
@@
Figure 7: Eigenvectors of edge states in the 3-layer case ($(u,\gamma)=(2,1)$) at $E=0$. (a) is the eigenvector as a function of $y$ at $(k_{x},E)=(0,0)$ and (b) at $(k_{x},E)=(-1,0)$. The edge is located within 1<y<1-1<y<1. $A_{n}$ and $B_{n}$ labels correspond to the $A$ and $B$ sub-lattice components of layer $n$ following the notation of [38]. Polarization is seen biased toward the outer layers.
One particularly interesting feature shown in Figure 4 and 5 is the flattening of the bulk spectrum above and below the band gap. This feature can in fact be straightforwardly deduced from the rotational invariance of the bulk Hamiltonian. Note that Figure 2 shows just a cross-section in $|\mathbf{k}|$ of the bulk spectrum. First, fix any $E$ which is not in the bulk band gap, and denote $k_{max}=\arg\max_{|\mathbf{k}|}\{E_{j}(\mathbf{k})=E,j\in\{1,...,2m\}\}$. Then clearly for any $k_{x}$ with $|k_{x}|\leq k_{max}$ we can find a $\mathbf{k}$ with $|\mathbf{k}|=\sqrt{k_{x}^{2}+k_{y}^{2}}=k_{max}$ for some $k_{y}\in[0,k_{max}]$ so, due to the rotation invariance of the spectrum, there is in fact a bulk mode corresponding to any $k_{x}\leq k_{max}$ for each $E$ not in the bulk band gap. Therefore the bulk spectrum of $H_{I}(k_{x})$ must be flat in between any two peaks in the bulk spectrum, as we observe in Figures 4, 5, 6, and 8.
@@ 5 Discussion
The result of Theorem 2.3 gives all possible topological values of the quantum anomalous Hall effect in an idealized macroscopic (effective) model of multi-layer rhombohedral graphene and Floquet topological insulators in a setting where the gate potential varies between two bulk regions. In the latter application, we retrieve the topological invariant obtained for large values of the driving laser frequency (corresponding to a small coupling constant $\gamma$) [3]. In the former application, we retrieve the topological invariant obtained for small displacement field [21, 36] compared to the coupling constant $\gamma$.
@@
We also note that the spatially varying gate potential is just one possible way to introduce non-trivial topological phases. One other possibility is the case that across a large enough length scale the energetically preferable stacking order changes continuously from ABC to CBA, as considered in [2]. This transition is modeled by a continuous transition of $B\to B^{*}$. Substituting the eigenvectors of $\hat{H}(u,\gamma A^{*})$ for $\hat{H}(-u,\gamma A)$ in Section 3 using the unitary relations derived in (8) it is clear that the BDI’s for this transition are identical a symmetric transition from phase j-j to $j$ as derived in Section 3. Figure 8 shows for the 5-layer case that indeed these two transitions are topologically equivalent. Indeed considering that $B$ and $B^{*}$ are distinct topological phases for each value of $u$ the number of topological phases allowing both the transition matrix $B$ and gate potential $u$ to vary spatially is $4\lfloor\frac{m}{2}\rfloor$. Bulk invariants for all these phases may be calculated similarly to in section 3.d and subsequently subtracting any two of these will give a BDI for a transition between any two of the $4\lfloor\frac{m}{2}\rfloor$ phases. One could also allow the possibility that a non-uniform electric field produces potential differences that are not constant between any two adjacent layers or adding insulating layers between graphene layers to introduce non-uniform coupling constants $\gamma$, which is the subject of ongoing work.
@@
Figure 8: Edge spectrum for 5-layer case with $A\to A^{*}$ transition. Parameter values used are $u=(4,1.2)$ for (a), and (b) respectively and $\gamma=1$. Comparison with Figure 4 $(m=5)$ shows that the $A\to A^{*}$ transition is topologically equivalent to the $u\to-u$ transition.
In the RHG application, we assume here a model (1) that is both spin- and valley-polarized. While the obtained numbers do not depend on spin, they depend on the valley index $\tau=\pm 1$. The total QAH therefore vanishes unless valley polarization may be obtained experimentally, which is achieved by means of an appropriate spin orbit coupling[21]. It is unclear whether such polarization may still be achieved for the large displacement fields (large values of $u$) necessary to observe a phase transition (i.e., $\delta<\delta_{\lfloor\frac{m}{2}-1\rfloor}$).
The theoretical results, and in particular the presence of a gap at $E=0$, depend on the specific structures of $H$ in (1). More general interlayer couplings such as those proposed in [21, 36] do not modify the elliptic nature of the operators and the bulk-edge correspondence still applies[7, 30]. Therefore, any continuous deformation of the model in (1) with bulk phases that does not close the gap at $E=0$ will generate an edge asymmetry for which the results and explicit indices of Theorem 2.3 apply. While the results of Theorem 2.2 guarantee that gap crossings may only occur at $E=0$, we found examples of values of $(\gamma,u)$ such that the spectral gap at values $k\not=0$ may be significantly smaller than the gap at $k=0$; see for instance Fig. 3. When considering interactions between more distant layers, as in [38], rotational symmetry which has been essential to our analysis here breaks down and the perturbative effects of further off-diagonal elements have yet to be explored. It is therefore unclear whether such gaps persist except for very small perturbations of the model (1).
@@
In summary, we have found all the topological phases and bulk difference invariants of an effective model of coupled 2-d Dirac systems which has applications in gated ABC-stacked rhombohedral graphene with an arbitrary number of layers [38, 27, 36] and Floquet topological insulators [3]. Due to the ellipticity of the problem the bulk-edge correspondence holds [7, 6] and we verify by numerical spectral calculations the number of predicted edge modes for all topological phases in the $m=\{3,4,5,6\}$ cases. For $u<\gamma$ we recover the results of [36] for the QAH effect in RHG and for $u\gg\gamma$ the results of [3] for Floquet topological insulators. Our results show that for an $m$-layer model with a spatially varying gate potential there are in fact $2\lfloor\frac{m}{2}\rfloor$ topological phases and the asymmetric charge transport at an interface between any two of these phases is quantized by the BDI in (15). Although it remains to be seen if these phases are robust to perturbations in interactions beyond the nearest-layer interactions we consider here, these results show the exciting possibility of tunable QAH states in multi-layer ABC stacked graphene should experimental conditions allow high enough gate potentials to be achieved.
Acknowledgments
The authors acknowledge stimulating discussions with Allan MacDonald on graphene systems and Jeremy Hoskins for invaluable advice and input regarding numerical techniques. This work was funded in part by NSF grant DMS-230641 and ONR grant N00014-26-1-2017.
Author Declarations
The authors have no conflicts to disclose.
Data Availability
The code needed to reproduce the data in section 4 is available at:
https://doi.org/10.5281/zenodo.17156250.
Appendix A Classification of elliptic Hamiltonians